Find an equation of the line that contains the following pairs of points? (3,3) (1,6)
Answer by Alan3354(69443)

Find an equation of the line that contains the following pairs of points? (3,3) (1,6)
---------------------
Do it like this:
i need to find the equation of the line that passes through each pai of points and write the answer in standard form
(-2,-4), (1,-1)
--------------
Find the slope, m
m = diffy/diffx = (-4 + 1)/(-2 -1) = -3/-3 = 1
-----
Use y = mx + b and either point to find b
-1 = 1*1 + b
b = -2
-----
y = x - 2
x - y = 2
==================
Using determinants:
|+x +y 1|
|-2 -4 1| = 0
|+1 -1 1|
-------
x*(-4+1) - y*(-2-1) + 1*(2+4) = 0
-3x + 3y = -6
x - y = 2